# TW-CRL: Time-Weighted Contrastive Reward Learning for Efficient Inverse Reinforcement Learning

### v1

@@ Abstract
Episodic tasks in Reinforcement Learning (RL) often pose challenges due to sparse reward signals and high-dimensional state spaces, which hinder efficient learning. Additionally, these tasks often feature hidden “trap states”—irreversible failures that prevent task completion but do not provide explicit negative rewards to guide agents away from repeated errors. To address these issues, we propose Time-Weighted Contrastive Reward Learning (TW-CRL), an Inverse Reinforcement Learning (IRL) framework that leverages both successful and failed demonstrations. By incorporating temporal information, TW-CRL learns a dense reward function that identifies critical states associated with success or failure. This approach not only enables agents to avoid trap states but also encourages meaningful exploration beyond simple imitation of expert trajectories. Empirical evaluations on navigation tasks and robotic manipulation benchmarks demonstrate that TW-CRL surpasses state-of-the-art methods, achieving improved efficiency and robustness.
Machine Learning, ICML
@@ 1 Introduction
An episodic task is a common Reinforcement Learning (RL) problem that has a well-defined termination condition (Sutton et al., 1999). Many real-world problems, such as robotic manipulation and navigation, are episodic by nature, often featuring complex state and action spaces alongside sparse reward signals (Andrychowicz et al., 2018; Zhu et al., 2020; Yu et al., 2019). The sparsity of rewards significantly complicates the training process by providing a positive reward only when the goal is achieved, leaving most steps uninformative. As a result, agents may spend excessive time in states that provide no meaningful feedback, slowing down the learning process. Moreover, certain “trap states” can make the goal permanently unreachable after a single mistake. These states do not explicitly offer negative rewards, but act as hidden traps that derail the agent’s attempts to succeed.
@@
Designing dense reward functions is an effective approach to address challenges in RL environments with sparse rewards (Gehring et al., 2022; Sutton et al., 2011; Gershman & Daw, 2017; Chan et al., 2024). However, how to develop such reward functions remains a challenging task. A straightforward approach is to use distance-based rewards, where the agent receives rewards based on its distance from the goal (Trott et al., 2019). While this method is intuitive, distance-based reward may not be effective under more complex environments where the distance between current state and goal state is difficult to measure. Another promising approach to providing dense rewards is Inverse Reinforcement Learning (IRL) (Ng & Russell, 2000; Abbeel & Ng, 2010), which infer reward functions from expert demonstrations. These methods assign higher rewards to states or state-action pairs that more closely align with expert behavior; however, this imitation-based strategy can limit exploration and cause agents to converge to suboptimal behaviors (Mavor-Parker et al., 2022). As a result, agents may fail to discover alternative strategies, such as more efficient paths or ways to avoid trap states (Section 4.1). Additionally, most IRL-based approaches rely solely on successful demonstrations, which reduces data efficiency, especially in complex environments where failures are common during early training (Shiarlis et al., 2016). By ignoring failed experiences, these methods miss valuable information that could help agents learn to avoid undesirable trajectories, ultimately slowing down the learning process and reducing the robustness of the learned policies.
@@
To address these challenges, we propose Time-Weighted Contrastive Reward Learning (TW-CRL), a novel IRL approach that trains the agent based on a learned accurate and environment-aligned dense reward function. TW-CRL contains two key components: Time-Weighted function and Contrastive Reward Learning. The Time-Weighted function incorporates temporal information from trajectories, providing denser and more informative feedback. This allows the agent to receive more accurate rewards that reflect the importance of different states over time. Contrastive Reward Learning utilizes both successful and failed demonstrations, which enables the agent to efficiently locate goals while avoiding repeatedly falling into the same failure patterns, ultimately improving learning efficiency and task success rates.
@@ 2 Related Work
IRL Approaches Most existing IRL methods focus primarily on imitating expert behavior. Maximum Entropy IRL (MaxEnt) (Ziebart et al., 2008) encourages diverse actions while maintaining consistency with expert demonstrations by maximizing entropy. Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) applies adversarial learning to align the agent’s state-action distribution with that of the expert, while Adversarial Inverse Reinforcement Learning (AIRL) (Fu et al., 2017) improves upon this by decoupling the reward function from environment dynamics, enhancing generalization. Although recent research (Hugessen et al., 2024; Naranjo-Campos et al., 2024; Liu & Zhu, 2025) continues to advance these methods, they often overlook the valuable insights that can be gained from failed demonstrations. As a result, their ability to recognize and avoid potential traps in complex environments is limited. TW-CRL overcomes this challenge by leveraging both successful and failed demonstrations, allowing agents to explore beyond simple imitation.
@@
Learning from Failures in RL Recent studies have explored leveraging failures to improve learning. Inverse Reinforcement Learning from Failure (IRLF) (Shiarlis et al., 2016) helps agents imitate expert behavior while avoiding past failures but does not fully analyze failure patterns. Self-Adaptive Reward Shaping (SASR) (Ma et al., 2024) adjusts rewards based on past success rates to balance exploration and exploitation. Other works (Gao et al., 2019; Xie et al., 2019; Guo et al., 2023; Wu et al., 2024) also focus on failure utilization but overlook the timing and sequence of failures. Unlike these methods, TW-CRL incorporates temporal information from both successful and failed demonstrations, leading to a more informative reward function.
@@
Risk Awareness in RL Risk-aware methods in RL focus on training agents to avoid risky situations (García & Fernández, 2015). Approaches such as Risk-Averse Imitation Learning (RAIL) (Santara et al., 2017; Lam et al., 2022; Jaimungal et al., 2021) and Risk-Sensitive Generative Adversarial Imitation Learning (RS-GAIL) (Lacotte et al., 2018) use the Conditional Value at Risk (CVaR) metric to assess and manage risk, while Inverse Risk-Sensitive Reinforcement Learning (IRSRL) (Ratliff & Mazumdar, 2017) employs convex risk measures to minimize potential dangers. Recent studies (Fei et al., 2021; Zhao et al., 2025; Yang et al., 2024; Sun et al., 2023) also follow similar strategies by relying on predefined risk measures based on expert knowledge or historical data. In contrast, TW-CRL automatically identifies and avoids risky regions without requiring prior knowledge or explicit risk metrics.
3 Preliminary
@@
Episodic MDP: An episodic MDP (Sutton et al., 1999; Domingues et al., 2020; Neu & Pike-Burke, 2020) is formally defined as a tuple $({\mathcal{S}},\mathcal{A},T,\mu,P,r)$, where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space, $T$ is the number of steps in one episode, $\mu$ is the distribution of the initial state, $P:{\mathcal{S}}\times\mathcal{A}\rightarrow{\mathcal{S}}$ is the dynamics which gives the distribution of the next state $s^{\prime}$ given the current state $s$ and the action $a$, and $r$ is the reward function given by the environment. Episodic tasks, modeled by episodic MDPs, have a well-defined termination condition, meaning the agent eventually reaches a terminal state.
Inverse Reinforcement Learning: IRL (Ng & Russell, 2000; Abbeel & Ng, 2004) is a specialized approach within Imitation Learning (IL) that aims to infer the underlying reward function from expert demonstrations, rather than relying on a predefined reward signal provided by the environment. Unlike other IL algorithms like Behaviour Cloning (BC) (Bain & Sammut, 1995), which directly learn a policy from demonstrations, IRL focuses on uncovering the expert’s intent by analyzing the demonstrated trajectories (Arora & Doshi, 2020a).
The primary objective of IRL is to recover a reward function that explains the observed behavior of an expert (Metelli et al., 2017). This approach is particularly valuable in scenarios where specifying an appropriate reward function is challenging or when the goal is to understand the underlying motivations of an agent’s behavior (Arora & Doshi, 2020b). By learning the reward function, IRL methods aim to capture the essence of the task, potentially allowing for better generalization to new situations and providing more interpretable results compared to direct policy imitation.
Figure 1: Overview of TW-CRL. TW-CRL processes demonstrations, including expert demonstrations (containing only successful trajectories) and trajectories collected through agent-environment interactions (containing both successful and failed trajectories). In stage (1), labels ( $r_{\text{est}}$ ) are assigned to states in the trajectories using the Time-Weighted function. In stage (2), the reward function is updated with Contrastive Reward Learning loss. The updated reward function is then used to provide rewards to the agent, which, along with the state from the environment, helps train the policy and generate actions.
@@ 4 Method
In this section, we introduce TW-CRL, a novel IRL method that utilizes temporal information from both failed and successful demonstrations in episodic tasks. In Section 4.1, we define the concept of trap states and discuss their impact on agent training within episodic environments. In Section 4.2, we describe how the Time-Weighted function extracts temporal information and detail the use of this information with Contrastive Reward Learning loss function to train a reward function that helps the agent identify unseen trap states and goal states. Finally, we outline the entire TW-CRL process.
4.1 The Trap State Problem
We first define the concept of trap states. In episodic tasks, the agent may reach a state where it is impossible to achieve the goal state, no matter what actions it takes. For example, in an object-pushing task, if the robot accidentally pushes an object off the table, it cannot recover and complete the task. These states are typically not explicitly defined or labeled in the environment and do not provide negative rewards to signal the agent’s mistakes. They act as unseen traps that impede the agent’s progress. We aim to create a reward function that gives negative feedback for these hidden trap states, helping the agent learn to avoid them during training. Concretely, we define the trap states as follows:
Definition 4.1 (Trap state).
A trap state is a state $s_{\text{trap}}\in{\mathcal{S}}$ such that, if $s_{\text{trap}}$ is a trap state, then regardless of action $a$, the next state $s_{\text{trap}}^{\prime}$ is always a trap state. Formally,
$\mathcal{S}_{\text{trap}}=\{s_{\text{trap}}\in{\mathcal{S}}\mid\sum\limits_{s^{\prime}\in\mathcal{S}_{\text{trap}}}P(s^{\prime}\mid s_{\text{trap}},a)=1,\forall a\in\mathcal{A}\}.$
Here, $s_{\text{trap}}$ is a trap state, $\mathcal{S}_{\text{trap}}$ is the set of all trap states, $a$ is any action in the action space $\mathcal{A}$ , and $s^{\prime}$ is the next state determined by the environment based on $s_{\text{trap}}$. This means once the agent enters such a state, it remains stuck, as every action it takes keeps it in the same situation.
Generalization of Trap States We then extend the definition of trap states to a broader situation, which applies to most cases—any situation where the agent fails to achieve the goal. Episodic tasks often have sparse rewards, meaning the agent only receives a positive reward upon reaching the goal and does not get helpful feedback during the rest of the process. As a result, the agent may spend a lot of time in states where it receives no meaningful information, making it harder to learn and improve, eventually leading to failure. Just like we want to give negative feedback for trap states, we also want to provide negative feedback for these uninformative states that contribute to failure. This can help the agent avoid repeating the same mistakes in future exploration and learning, improving its overall training efficiency.
Based on this idea, we consider all failed trajectories as a gradual process in which the agent moves toward a trap state. In other words, we assume that every failed attempt eventually leads the agent to a trap state. Based on this assumption, we propose a method to generate a meaningful reward function that can help the agent learn more effectively.
4.2 Time-Weighted Contrastive Reward Learning
4.2.1 Time-Weighted
This section describes how TW-CRL extracts temporal information from demonstrations. Building on the concept of trap states discussed in Section 4.1, we assume that in a successful demonstration $\tau^{+}$, the final state is a goal state, while in a failed demonstration $\tau^{-}$, the final state is a trap state. Formally, in an episodic task, a successful demonstration $\tau^{+}$ is defined as a sequence of state-action pairs:
$\tau^{+}=\{(s_{1}^{+},a_{1}^{+}),(s_{2}^{+},a_{2}^{+}),\ldots,(s_{T}^{+},a_{T}^{+})\}.$
where $s_{T}^{+}\in\mathcal{S}_{\text{goal}}$. Let $\mathcal{D}^{+}=\{\tau^{+}|s_{T}^{+}\in\mathcal{S}_{\text{goal}}\}$ be the set of all successful demonstrations. In contrary, a failed demonstration $\tau^{-}$ is given by:
$\tau^{-}=\{(s_{1}^{-},a_{1}^{-}),(s_{2}^{-},a_{2}^{-}),\ldots,(s_{T}^{-},a_{T}^{-})\}.$
where $s_{T}^{-}\in\mathcal{S}_{\text{trap}}$. Let $\mathcal{D}^{-}=\{\tau^{-}|s_{T}^{-}\in\mathcal{S}_{\text{trap}}\}$ be the set of all failed demonstrations.
@@ 4.2.1 Time-Weighted
Based on the assumption above, in successful demonstrations, the agent gradually moves closer to the goal as the episode progresses, making states that appear later in the trajectory more likely to be goal states. Conversely, in failed demonstrations, the agent eventually reaches a trap state that prevents task completion, meaning that states appearing later in the trajectory are more likely to be trap states. Additionally, in episodic tasks, the first few states have a lower probability of being goal or trap states.
Therefore, we aim to use a concave function to estimate the probability that the current state is a goal or trap state as time $t$ increases. We introduce a Time-Weighted function $w(t)$ to represent the estimated probability that a state $s_{t}$ within a certain demonstration $\tau$ is a goal state ($P(s_{t}\in\mathcal{S}_{\text{goal}})$) or a trap state ($P(s_{t}\in\mathcal{S}_{\text{trap}})$) based on temporal information. As $t$ increases, the likelihood that $s_{t}$ is a trap state or a goal state also increases, and the growth rate becomes more pronounced with larger $t$.
To capture this temporal dynamic, we define the Time-Weighted function as follows:
$w(t)=\frac{t}{T}\exp\left(k\frac{t}{T}\right),$ (1)
where $t\in\{1,\ldots,T\}$ represents the timestep, $T$ is the episode horizon, and $k$ is a hyperparameter that controls the rate of weight growth. The exponential term $\exp(\cdot)$ ensures that the weight increases more significantly as the episode progresses, emphasizing the importance of later states. This helps TW-CRL to train a more effective reward function.
4.2.2 Contrastive Reward Learning

### v2

@@ Abstract
Episodic tasks in Reinforcement Learning (RL) often pose challenges due to sparse reward signals and high-dimensional state spaces, which hinder efficient learning. Additionally, these tasks often feature hidden “trap states”—irreversible failures that prevent task completion but do not provide explicit negative rewards to guide agents away from repeated errors. To address these issues, we propose Time-Weighted Contrastive Reward Learning (TW-CRL), an Inverse Reinforcement Learning (IRL) framework that leverages both successful and failed demonstrations. By incorporating temporal information, TW-CRL learns a dense reward function that identifies critical states associated with success or failure. This approach not only enables agents to avoid trap states but also encourages meaningful exploration beyond simple imitation of expert trajectories. Empirical evaluations on eight benchmarks demonstrate that TW-CRL surpasses state-of-the-art methods, achieving improved efficiency and robustness.
@@ 1 Introduction
An episodic task is a common Reinforcement Learning (RL) problem that has a well-defined termination condition [35]. Many real-world problems, such as robotic manipulation and navigation, are episodic by nature, often featuring complex state and action spaces alongside sparse reward signals [2, 45, 43]. The sparsity of rewards significantly complicates the training process by providing a positive reward only when the goal is achieved, leaving most steps uninformative. As a result, agents may spend excessive time in states that provide no meaningful feedback, slowing down the learning process. Moreover, certain “trap states” can make the goal permanently unreachable after a single mistake. These states do not explicitly offer negative rewards, but act as hidden traps that derail the agent’s attempts to succeed.
@@
Designing dense reward functions is an effective approach to address challenges in RL environments with sparse rewards [14, 34, 15, 3]. However, how to develop such reward functions remains a challenging task. A straightforward approach is to use distance-based rewards, where the agent receives rewards based on its distance from the goal [38]. While this method is intuitive, distance-based reward may not be effective under more complex environments where the distance between current state and goal state is difficult to measure. Another promising approach to providing dense rewards is Inverse Reinforcement Learning (IRL) [27, 1], which infer reward functions from expert demonstrations. These methods assign higher rewards to states or state-action pairs that more closely align with expert behavior; however, this imitation-based strategy can limit exploration and cause agents to converge to suboptimal behaviors [24]. As a result, agents may fail to discover alternative strategies, such as more efficient paths or ways to avoid trap states (Section 3.2). Additionally, most IRL-based approaches rely solely on successful demonstrations, which reduces data efficiency, especially in complex environments where failures are common during early training [32]. By ignoring failed experiences, these methods miss valuable information that could help agents learn to avoid undesirable trajectories, ultimately slowing down the learning process and reducing the robustness of the learned policies.
@@
To address these challenges, we propose Time-Weighted Contrastive Reward Learning (TW-CRL), a novel IRL approach that trains the agent based on a learned accurate and environment-aligned dense reward function. TW-CRL contains two key components: Time-Weighted function and Contrastive Reward Learning. The Time-Weighted function incorporates temporal information from trajectories, providing denser and more informative feedback. This allows the agent to receive more accurate rewards that reflect the importance of different states over time. Contrastive Reward Learning utilizes both successful and failed demonstrations, which enables the agent to efficiently locate goals while avoiding repeatedly falling into the same failure patterns, ultimately improving learning efficiency and average episodic return.
@@ 2 Related Work
IRL Approaches Most existing IRL methods focus primarily on imitating expert behavior. Maximum Entropy IRL (MaxEnt) [46] encourages diverse actions while maintaining consistency with expert demonstrations by maximizing entropy. Generative Adversarial Imitation Learning (GAIL) [17] and Adversarial Inverse Reinforcement Learning (AIRL) [9] applies adversarial learning to align the agent’s state-action distribution with that of the expert. Although recent research [18, 25, 22] continues to advance these methods, they often overlook the valuable insights that can be gained from failed demonstrations. As a result, their ability to recognize and avoid potential traps in complex environments is limited. TW-CRL overcomes this challenge by leveraging both successful and failed demonstrations, allowing agents to explore beyond simple imitation.
@@
Learning from Failures in RL Recent studies have explored leveraging failures to improve learning. Inverse Reinforcement Learning from Failure (IRLF) [32] helps agents imitate expert behavior while avoiding past failures but does not fully analyze failure patterns. Self-Adaptive Reward Shaping (SASR) [23] adjusts rewards based on past success rates to balance exploration and exploitation. Other approaches such as [4, 30, 39] and [18] incorporate suboptimal demonstrations to aid training but typically rely only on successful yet imperfect trajectories, ignoring completely failed ones. [12, 41, 16, 40] also focus on failure utilization but overlook the timing and sequence of failures. Unlike these methods, TW-CRL incorporates temporal information from both successful and failed demonstrations, leading to a more informative reward function.
@@
Risk Awareness in RL Risk-aware methods in RL focus on training agents to avoid risky situations [13]. Approaches such as Risk-Averse Imitation Learning (RAIL) [31, 21, 19] and Risk-Sensitive Generative Adversarial Imitation Learning (RS-GAIL) [20] use the Conditional Value at Risk (CVaR) metric to assess and manage risk, while Inverse Risk-Sensitive Reinforcement Learning (IRSRL) [29] employs convex risk measures to minimize potential dangers. Recent studies [7, 44, 42, 33] also follow similar strategies by relying on predefined risk measures based on expert knowledge or historical data. In contrast, TW-CRL automatically identifies and avoids risky regions without requiring prior knowledge or explicit risk metrics.
3 Problem Setting
3.1 Episodic Tasks
@@
An episodic MDP [35, 6, 26] is formally defined as a tuple $({\mathcal{S}},\mathcal{A},T,\mu,P,r)$, where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space, $T$ is the number of steps in one episode, $\mu$ is the distribution of the initial state, $P:{\mathcal{S}}\times\mathcal{A}\rightarrow{\mathcal{S}}$ is the dynamics which gives the distribution of the next state $s^{\prime}$ given the current state $s$ and the action $a$, and $r$ is the reward function given by the environment. Episodic tasks, modeled by episodic MDPs, have a well-defined termination condition. In contrast, continuous tasks are tasks without clear termination conditions and proceeds indefinitely.
3.2 Trap States
Consider an episodic task with state space $\mathcal{S}$. For now, we assume that the task has a clear set of successful terminal states, meaning that there exists a non-empty set of goal states $\mathcal{S}_{\text{goal}}\subseteq\mathcal{S}$ that contains all terminal states indicating successful completion. Once the agent reaches a goal state $s_{\text{goal}}\in\mathcal{S}_{\text{goal}}$, it remains in the same state until the end of that episode. We’ll extend our problem setting to other task types in Section 4.4.
Given these assumptions, we define the concept of trap states. In episodic tasks, the agent may reach a state where it is impossible to reach a goal state, no matter what actions it takes. For example, in an object-pushing task, if the robot accidentally pushes an object off the table, it cannot recover and complete the task. These states are typically not explicitly defined or labeled in the environment and do not provide negative rewards to signal the agent’s mistakes. They act as unseen traps that impede the agent’s progress. We aim to create a reward function that gives negative feedback for these hidden trap states, helping the agent learn to avoid them during training. Concretely, we define the trap states as follows:
Definition 3.1 (Trap state).
A trap state is a state $s_{\text{trap}}\in{\mathcal{S}}$ such that, if $s_{\text{trap}}$ is a trap state, then regardless of action $a$, the next state $s_{\text{trap}}^{\prime}$ is always a trap state. Additionally, all goal states are not trap states. Formally,
$\mathcal{S}_{\text{trap}}=\{s_{\text{trap}}\in{\mathcal{S}}-\mathcal{S}_{\text{goal}}\mid\sum\limits_{s^{\prime}\in\mathcal{S}_{\text{trap}}}P(s^{\prime}\mid s_{\text{trap}},a)=1,\forall a\in\mathcal{A}\}.$ (1)
Here, $s_{\text{trap}}$ is a trap state, $\mathcal{S}_{\text{trap}}$ is the set of all trap states, $a$ is any action in the action space $\mathcal{A}$ , and $s^{\prime}$ is the next state determined by the environment based on $s_{\text{trap}}$. This means once the agent enters such a state, it remains stuck in a set of trap states, as every action it takes keeps it in the same situation.
3.3 Generalization of Trap States
We then extend the definition of trap states to a broader situation, which applies to most cases—any situation where the agent fails to achieve the goal. Episodic tasks often have sparse rewards, meaning the agent only receives a positive reward upon reaching the goal and does not get helpful feedback during the rest of the process. As a result, the agent may spend a lot of time in states where it receives no meaningful information, making it harder to learn and improve, eventually leading to failure. Just like we want to give negative feedback for trap states, we also want to provide negative feedback for these uninformative states that contribute to failure. This can help the agent avoid repeating the same mistakes in future exploration and learning, improving its overall training efficiency.
Based on this idea, we consider all failed trajectories as a gradual process in which the agent moves toward a trap state. In other words, we assume that the final state of every failed trajectory is a trap state. Formally, in an episodic task, a failed trajectory $\tau^{-}$ is defined as a sequence of state-action pairs:
$\tau^{-}=\{(s_{1}^{-},a_{1}^{-}),(s_{2}^{-},a_{2}^{-}),\ldots,(s_{T}^{-},a_{T}^{-})\}.$ (2)
where $s_{T}^{-}\in\mathcal{S}_{\text{trap}}$. Contrarily, in an episodic task, the final state of a successful trajectory is always a goal state. Therefore, a successful trajectory $\tau^{+}$ can be represented similarly as:
$\tau^{+}=\{(s_{1}^{+},a_{1}^{+}),(s_{2}^{+},a_{2}^{+}),\ldots,(s_{T}^{+},a_{T}^{+})\}.$ (3)
where $s_{T}^{+}\in\mathcal{S}_{\text{goal}}$. Furthermore, we can apply the notion above to a set of expert demonstrations $\mathcal{D}$. We denote the set of all successful demonstrations as $\mathcal{D}^{+}=\{\tau^{+}\in\mathcal{D}|s_{T}^{+}\in\mathcal{S}_{\text{goal}}\}$. The set of all failed demonstrations is denoted as $\mathcal{D}^{-}=\{\tau^{-}\in\mathcal{D}|s_{T}^{-}\in\mathcal{S}_{\text{trap}}\}$.
Based on these assumptions, we propose a method to generate a meaningful reward function that can help the agent learn more effectively.
@@ 4 Method
In this section, we introduce TW-CRL, a novel IRL method that utilizes temporal information from both failed and successful demonstrations in episodic tasks. We first describe how we use the Time-Weighted function to extract temporal information from trajectories (Section 4.1). Next, we detail the use of Contrastive Reward Learning loss function to train a reward function that helps the agent identify unseen trap states and goal states (Section 4.2). Then, we outline the entire TW-CRL process (Section 4.3). Finally, we discuss how TW-CRL could be extended to tasks types beyond episodic tasks with clear sets of successful terminal states (Section 4.4).
Figure 1: Overview of TW-CRL.
4.1 Time-Weighted Function
@@
We first describe how TW-CRL extracts temporal information from demonstrations. As discussed in Section 3.2, in successful demonstrations, the agent gradually moves closer to the goal as the episode progresses, making states that appear later in the trajectory more likely to be goal states. Conversely, in failed demonstrations, the agent eventually reaches a trap state that prevents task completion, meaning that states appearing later in the trajectory are more likely to be trap states. To capture this temporal feature, we model the dynamics of the agent as an absorbing Markov chain. For failed demonstrations, the trap states are absorbing states and non-trap states are transient states. In a failed demonstration, recall that we assume the final state is always a trap state ($s_{T}\in\mathcal{S}_{\text{trap}}$). Additionally, we also assume that the initial state is not a trap state ($s_{0}\notin\mathcal{S}_{\text{trap}}$). Under these assumptions, given the prior knowledge that a demonstration is a failed demonstration, we can derive the conditional probability of a state $s_{t}$ of timestep $t$ being a trap state as follows:
$P(s_{t}\in\mathcal{S}_{\text{trap}}|s_{T}\in\mathcal{S}_{\text{trap}})$ $=\frac{1-(1-f(t,T))^{t}}{1-(1-f(t,T))^{T}}$ (4)
where $f:\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{R}$ is a function that indicates the probability of a non-trap state transitioning into a trap state in timestep $t$ given total horizon $T$. The derivation of Equation 4 is in Appendix A.1. In practice, $f(t,T)$ can be adjusted to meet different needs of specific environments. In our experiments, we let $f(t,T)\propto e^{\alpha t}$, where $\alpha>0$ is a hyperparameter. This is because we observe an exponential increase with respect to the timestep matches the best with our environments. After normalization to ensure that $f(0,T)=0$ and $f(T,T)=1$, we have
$f(t,T)=\frac{e^{\alpha t}-1}{e^{\alpha T}-1}$ (5)
Given this assumption, Equation 4 can be further simplified:
$P(s_{t}\in\mathcal{S}_{\text{trap}}|s_{T}\in\mathcal{S}_{\text{trap}})\approx 1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}$ (6)
For successful demonstrations, the goal states are absorbing states, and non-goal states are transient states. Similar to our conclusion above, we can derive that:
$P(s_{t}\in\mathcal{S}_{\text{goal}}|s_{T}\in\mathcal{S}_{\text{goal}})\approx 1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}$ (7)
The derivations of Equation 6 and (7) are in Appendix A.2. We can observe that the right hand side of Equation 6 and (7) are the same. We thus use it as our Time-Weighted function $w(t)$ for both successful and failed demonstrations:
$w(t)=1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}$ (8)
Further discussions regarding the properties of $w(t)$ is in Appendix A.3. As we’ll discuss in Section 4.2, the Time-Weighted function is used as a weighting function for states in different timesteps within each demonstration. By adopting the simplified conditional probabilities derived above as our Time-Weighted function, we make sure that higher weights are assigned to states that have heavier impacts on the success or failure of the trajectory they’re in. Specifically, the weight assigned to each state equals the probability of the agent entering $\mathcal{S}_{\text{goal}}$ / $\mathcal{S}_{\text{trap}}$ in that timestep.
4.2 Contrastive Reward Learning